Grade 11Mechanics

how does the relation between escape velocity and orbital velocity came, derive the relation??

pratik
10 Years agoGrade 11

1 Answer

Yash verma
10 Years ago
Ve2=2gr
Vo2=gr
Ve2=2Vo2
Hence,                Ve=21/2 Vo  (Here r=Re +h ; if h
 
Orbital velocity is the speed that an object must maintain in order to orbit another object. Escape velocity is the speed at which an object would break orbit. Escape velocity must be square-root of 2 times larger than the orbital velocity in order to break orbit. (l) If the velocities are equal, the object will stay in a constant orbit, not changing in elevation. (ll) less than orbital the orbit will decay and the object will crash. (lll) more than orbital and the object will have an accending orbit and fly off into space.